ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

4. Algebraic Geometry

1991 MS Classification: 81T30 Dijkgraaf, Robbert, University of Amsterdam The Mathematics of Fivebranes (See section 11)

ICM 1998 Berlin · Plenary and Invited Lectures